OpenStudy (anonymous):

Find the indefinite integral of y = (sqrt(9x^(2/3) + 4))/(3(cbrt(x)))

OpenStudy (dumbcow):

use substitution let \[u = 9x^{2/3} +4\] then \[du = \frac{6}{x^{1/3}} dx\] changing the integral to \[\int\limits_{}^{}2 \sqrt{u} du\]

OpenStudy (dumbcow):

oops sorry the last part should be \[\int\limits_{}^{}\frac{\sqrt{u}}{18} du\]